OpenStudy (kj4uts):

The function f(x)=2x^3-4x+1 has the output value 12 for exactly one input value of x. Between which of the following pairs of numbers does this input value of x lie? 2.08 and 2.09 2.10 and 2.11 2.12 and 2.13 2.14 and 2.15 2.16 and 2.17 2.19 and 2.20 None of the above Please explain. Thank you!

OpenStudy (kj4uts):

How do I do this do I plug the numbers in?

hero (hero):

By plugging them in for x

OpenStudy (kj4uts):

So plug in for example 2(2.08)^3-4(2.08)+1=10.68 and then 2(2.09)^3-4(2.09)+1=10.89

hero (hero):

Keep trying until you get close enough to the exact value.

OpenStudy (kj4uts):

for 2(2.14)^3-4(2.14)+1 I got 12.04 and 2(2.15)^3-4(2.15)+1 I got 12.28

hero (hero):

Keep trying until you get close enough without going over.

OpenStudy (kj4uts):

and then for 2.16 and 2.17 I got 12.51 - 12.76

OpenStudy (kj4uts):

@hero so what would it be?

OpenStudy (kj4uts):

@hero or is it none of the above

hero (hero):

Why'd you skip 2.09 thru 2.13

OpenStudy (kj4uts):

@hero I did those but those were closer to 12 2.08 and 2.09 = 10.68 - 10.89 2.10 and 2.11 = 11.12 - 11.81 2.12 and 2.13 = 11.58 - 11.81

hero (hero):

Try 2.13-2.14

hero (hero):

And that should let you know what the answer is

OpenStudy (kj4uts):

but that's not a choice

OpenStudy (kj4uts):

@hero 11.81 - 12.04 i seen someone mention online that "The easiest way to do this is probably to try each pair of numbers and see which one gives numbers on either side of 12." which is what we have been doing but what does it mean by either side of 12?

hero (hero):

"above or below" 12

hero (hero):

on a number line

OpenStudy (kj4uts):

above and below 12 would be all of them except 2.14 and 2.15 and 2.16 and 2.17

hero (hero):

Not really.

OpenStudy (kj4uts):

@hero it doesn't look like none of them are the answer because 12 does not fall in between any of these numbers

hero (hero):

Okay.